On a map of the United States, the distance between Seattle and Durango is 4.125 in. The two cities are actually 1238 miles apart. On this same map, what would be the distance between Chicago and El Paso, two cities that are actually 1606 mi apart?
Answer by rfer(16322)

4.125/1238=x/1606
1238x=1606*4.125
x=6624.75/1238
x=5.35 in.